# Open-Source Reinforcement Learning Environments Implemented in MuJoCo with Franka Manipulator

### v1

@@ Abstract
Index Terms:
Reinforcement Learning, MuJoCo, Franka Emika Panda
@@ I INTRODUCTION
Issac Gym [16], built on NVIDIA PhysX, features accelerating the training process on GPU and also incorporates a host of robotics-related environments. Factory [17] resorts to reproduce the realistic contact-rich assembly scenario in simulator, and thus three environments are developed to solve the assembly tasks on the NIST task board, e.g., FrankaNutBoltEnv, FrankaInsertionEnv, and FrankaGearsEnv. Single and dual KUKA arm base environments with each of the three task variants (reorientation, regrasping, and grasp-and-throw) are also created in Issac Gym [18].
@@ II-C Truncated Quantile Critics
TQC is an extension of SAC but learn the distributional representation for $N$ Q-function critics. Each critic $Z_{\phi_{n}}(s_{t},a_{t}),n\in\{1,...,N\}$ estimates the distribution over the sum of discount rewards and contains $M$ atoms, i.e., $Z_{\phi_{n}}(s_{t},a_{t})=$ $\tfrac{1}{M}\sum_{m=1}^{M}\delta(\theta_{\phi_{n}}^{M}(s_{t},a_{t}))$. The temperal difference target distribution $y_{i}(s_{t},a_{t}):$ $=r(s_{t},a_{t})+\gamma[\theta_{\phi_{n}}^{m}(s_{t+1},a_{t+1})-\alpha\log\pi_{\theta}(\tilde{a}_{t+1}|s_{t+1})],m\in\{1,...,M\}$ is used to learn the quantile locations for fractions $\tau_{m}=\tfrac{2m-1}{M}$ with Huber quantile loss and then pool all atoms in one set. After sorting them in ascending order, $d\times N$ atoms are removed to control the overestimation error, and the remaining $k\times N$ smallest atoms are used to define the target distribution $Y(s_{t},a_{t})=\tfrac{1}{kN}\sum_{i=1}^{kN}\delta(y_{i}(s_{t},a_{t}))$. The parameters of approximators can be trained by minimizing
@@ III ENVIRONMENTS
Number of critics $2$ (TQC only)
@@
Quantiles to drop per critic $2$ (TQC only)
@@ IV EVALUATION
This section presents some training details and results using DDPG, SAC, and TQC, which are implemented with Stable-Baselines3[23]. At the beginning of each epiosde, the arm is reset to the initial configuration, and the positions of target and goal are randomized within the specified area. For FrankaPickAndPlace, the trick of the target position being randomized on the floor facilitates convergence, and the corresponding random probability can be adjusted. The episode length is 50 by default. Truncated and terminated signals are detected after each step. Termination can end the current episode in advance without wandering up to the maximum step limit. The transitions are saved in the replay buffer with Hindsight Experience Replay (HER) [24]. The policy is evaluated via 15 epiosdes at an interval of 2,000 training steps. The detailed hyperparameters are available in Table I.
Each algorithm is trained over 3 random seeds, and all training processes are deployed on a workstation with a NVIDIA RTX 3070Ti GPU and an i7-12700K CPU. Fig. 3 depicts the evaluation curves with the sparse reward for all tasks. For FrankaSlide, the target will be randomly distributed to the unreachable area for the Franka arm, which makes it hard to consistently adjust the object position. Thus, the training lasts 1 million steps to improve the agent’s performance. As shown in Fig. 2(b), SAC and TQC stabilize around the $80\%$ success rate after training and rarely increase further, which is similar to the distributions in panda-gym and Fetch. Besides, SAC and TQC perform well across all tasks and achieve a success rate close to 1 after training. Especially for FrankaPush, SAC and TQC reach a success rate of more than $80\%$ around $2\times 10^{5}$ steps. Finally, all tasks are completed successfully with the trained policies, which are shown in Fig. 4.

### v2

@@ I INTRODUCTION
Isaac Gym [16], built on NVIDIA PhysX, features accelerating the training process on GPU and also incorporates a host of robotics-related environments. Factory [17] resorts to reproduce the realistic contact-rich assembly scenario in simulator, and thus three environments are developed to solve the assembly tasks on the NIST task board, e.g., FrankaNutBoltEnv, FrankaInsertionEnv, and FrankaGearsEnv. Single and dual KUKA arm base environments with each of the three task variants (reorientation, regrasping, and grasp-and-throw) are also created in Issac Gym [18].
@@ II-C Truncated Quantile Critics
TQC is an extension of SAC but learn the distributional representation for $N$ Q-function critics. Each critic $Z_{\phi_{n}}(s_{t},a_{t}),n\in\{1,...,N\}$ estimates the distribution over the sum of discount rewards and contains $M$ atoms, i.e., $Z_{\phi_{n}}(s_{t},a_{t})=$ $\tfrac{1}{M}\sum_{m=1}^{M}\delta(\theta_{\phi_{n}}^{M}(s_{t},a_{t}))$. The temperal difference target distribution $y_{i}(s_{t},a_{t}):$ $r(s_{t},a_{t})+\gamma[\theta_{\phi_{n}}^{m}(s_{t+1},a_{t+1})-\alpha\log\pi_{\theta}(\tilde{a}_{t+1}|s_{t+1})],m\in\{1,...,M\}$ is used to learn the quantile locations for fractions $\tau_{m}=\tfrac{2m-1}{M}$ with Huber quantile loss and then pool all atoms in one set. After sorting them in ascending order, $d\times N$ atoms are removed to control the overestimation error, and the remaining $k\times N$ smallest atoms are used to define the target distribution $Y(s_{t},a_{t})=\tfrac{1}{kN}\sum_{i=1}^{kN}\delta(y_{i}(s_{t},a_{t}))$. The parameters of approximators can be trained by minimizing
@@ III ENVIRONMENTS
Number of critics 22 (TQC only)
@@
Quantiles to drop per critic 22 (TQC only)
@@ IV EVALUATION
This section presents some training details and results using DDPG, SAC, and TQC, which are implemented with Stable-Baselines3[23]. At the beginning of each epiosde, the arm is reset to the initial configuration, and the positions of target and goal are randomized within the specified area. For FrankaPickAndPlace, the trick of the target position being randomized on the floor facilitates convergence, and the corresponding random probability can be adjusted. The episode length is 50 by default. Truncated and terminated signals are detected after each step. Termination can end the current episode in advance without wandering up to the maximum step limit. The transitions are saved in the replay buffer with Hindsight Experience Replay (HER) [24]. The policy is evaluated via 50 epiosdes at an interval of 2,000 training steps. The detailed hyperparameters are available in Table I.
Each algorithm is trained over 3 random seeds, and all training processes are deployed on a workstation with a NVIDIA RTX 3070Ti GPU and an i7-12700K CPU. Fig. 3 depicts the evaluation curves with the sparse reward for all tasks. For FrankaSlide, the target will be randomly distributed to the unreachable area for the Franka arm, which makes it hard to consistently adjust the object position. Thus, the training lasts 1 million steps to improve the agent’s performance. As shown in Fig. 3(b), SAC and TQC stabilize around the $80\%$ success rate after training and rarely increase further, which is similar to the distributions in Fetch [12] and panda-gym [15]. Besides, SAC and TQC perform well across all tasks and achieve a success rate close to 1 after training. Especially for FrankaPush, SAC and TQC reach a success rate of more than $80\%$ around $2\times 10^{5}$ steps. Finally, all tasks are completed successfully with the trained policies, which are shown in Fig. 4.